# FlowCast: Advancing Precipitation Nowcasting with Conditional Flow Matching

### v2

@@ Abstract
Radar-based precipitation nowcasting, the task of forecasting short-term precipitation fields from previous radar images, is a critical problem for flood risk management and decision-making. While deep learning has substantially advanced this field, two challenges remain fundamental: the uncertainty of atmospheric dynamics and the efficient modeling of high-dimensional data. Diffusion models have shown strong promise by producing sharp, reliable forecasts, but their iterative sampling process is computationally prohibitive for time-critical applications. We introduce FlowCast, the first end-to-end probabilistic model leveraging Conditional Flow Matching (CFM) as a direct noise-to-data generative framework for precipitation nowcasting. Unlike hybrid approaches, FlowCast learns a direct noise-to-data mapping in a compressed latent space, enabling rapid, high-fidelity sample generation. Our experiments demonstrate that FlowCast establishes a new state-of-the-art in probabilistic performance while also exceeding deterministic baselines in predictive accuracy. A direct comparison further reveals the CFM objective is both more accurate and significantly more efficient than a diffusion objective on the same architecture, maintaining high performance with significantly fewer sampling steps. This work positions CFM as a powerful and practical alternative for high-dimensional spatiotemporal forecasting.
@@ 1 Introduction
Accurate and timely short-term precipitation forecasts, or nowcasting, are of paramount importance due to their significant socio-economic impacts, such as issuing flood warnings and managing water resources. Precipitation nowcasting, as defined in this work, involves predicting a sequence of future radar images from historical observations for the immediate future up to a few hours (an2024deep). Traditional methods, like Eulerian and Lagrangian persistence (germann2002scale), rely on advecting the current precipitation field. However, their simplified physical assumptions limit their ability to capture the complex, non-linear dynamics of atmospheric processes, especially for rapidly evolving weather systems (prudden2020review).
Deep learning has introduced a paradigm shift in precipitation nowcasting. Deterministic models based on recurrent and transformer architectures learn complex spatiotemporal patterns directly from large volumes of radar data (prudden2020review; an2024deep). While these models outperform traditional methods, optimizing for metrics like Mean Squared Error (MSE) compels them to produce a single, best-guess forecast. This often results in overly smooth predictions at longer lead times, failing to capture the inherent uncertainty in precipitation evolution and underrepresenting high-impact weather events.
To address this, probabilistic generative models have become central to modern nowcasting, aiming to predict a distribution over many plausible futures. Diffusion models (ho2020denoising), in particular, have emerged as the state-of-the-art, producing sharp and reliable ensemble forecasts (gao2024prediff; leinonen2023latent; gong2024cascast). However, this performance comes at a steep price: their reliance on an iterative denoising process, often requiring hundreds of function evaluations for a single forecast, makes them computationally expensive. This high Number of Function Evaluations (NFE) poses a significant barrier to practical application in time-critical scenarios where rapid ensemble generation is crucial.
This work introduces FlowCast, a novel probabilistic nowcasting model built on Conditional Flow Matching (CFM) (lipman2023flow; tong2024improving), a powerful and efficient alternative designed for rapid sampling. While recent work has applied rectified flows for the deterministic refinement of blurry forecasts (fengperceptually), FlowCast is, to our knowledge, the first to successfully apply CFM as a full, noise-to-data generative model for this task. We demonstrate that FlowCast alleviates the tension between accuracy and efficiency, establishing a new state-of-the-art by exceeding the performance of leading diffusion models while offering a superior performance-cost trade-off.
@@
  • We establish a new state-of-the-art in both probabilistic performance and predictive accuracy on two diverse radar datasets, the benchmark SEVIR dataset (veillette2020sevir) and the local ARSO dataset.
@@ 2.1 Deterministic Nowcasting
Deep learning for precipitation nowcasting has evolved from RNN-based architectures to Transformer-based models. Early work includes ConvLSTM (shi2015convolutional), extending LSTMs with convolutions for spatiotemporal data, and the PredRNN family (wang2017predrnn; wang2022predrnn), which introduced a spatiotemporal memory flow for improved long-range dependency modeling. More recently, Transformer architectures like Earthformer (gao2022earthformer) and Earthfarseer (wu2024earthfarsser) have set new benchmarks by using attention to model complex global dynamics. A common limitation of deterministic models is that they produce overly smooth forecasts when trained with pixel-wise losses (e.g., MSE), as they average over possible futures.
@@ 2.2 Probabilistic Nowcasting
GANs (ravuri2021skilful) were an early approach for producing sharp forecasts but suffer from training instability. Diffusion models (ho2020denoising) have recently emerged as the state-of-the-art, offering stable training and high-quality samples. PreDiff (gao2024prediff) and LDCast (leinonen2023latent) are prominent latent diffusion models for ensemble forecasting. A notable hybrid is CasCast (gong2024cascast), which uses a deterministic model for large-scale patterns and a conditional diffusion model to refine stochastic details.
@@
Generative modeling with flows offers an attractive alternative to diffusion. Traditional Continuous Normalizing Flows (CNFs) (chen2018neural) model data via ODEs but require expensive numerical integration during training to compute likelihoods, making them computationally prohibitive for high-dimensional spatiotemporal data. Conditional Flow Matching (CFM) (lipman2023flow) and Rectified Flows (liu2023flow) overcome this by regressing a vector field against a conditional probability path, enabling simulation-free training. However, while standard Rectified Flows typically utilize a singular conditional path (effectively $\sigma\to 0$), our application of Independent CFM (I-CFM) incorporates a Gaussian probability path with $\sigma>0$. This ”thickens” the training trajectory, providing crucial regularization that stabilizes the learning of the vector field for high-dimensional data compared to the singular paths of rectified flows.
Crucially, while diffusion models rely on stochastic denoising paths that are often curved and require many sampling steps, CFM allows for learning straight-line ODE trajectories between noise and data (tong2024improving). This geometric property enforces a direct mapping that preserves temporal coherence and allows for rapid sampling. While fengperceptually recently used a rectified flow module to strictly refine deterministic forecasts, FlowCast applies CFM as a standalone probabilistic generative model. This allows it to learn the full noise-to-data distribution and capture multimodal uncertainty without relying on a deterministic base forecast.
@@ 3.2 Latent Conditional Flow Matching
To reduce the high computational cost of generative modeling, we adopt a two-stage approach inspired by latent diffusion models (rombach2022high). A Variational Autoencoder (VAE) (kingma2013auto) compresses high-dimensional radar frames into low-dimensional latents, which are used to train a generative model in the latent space.
Our generative model is built on the Conditional Flow Matching (CFM) framework (lipman2023flow), which trains a continuous normalizing flow by learning a vector field $v_{\theta}$ that maps samples from a prior distribution (e.g., Gaussian) to the target data distribution. We use Independent CFM (I-CFM) (tong2024improving), which defines a probability path $p_{t}(x_{t}|x_{0},x_{1})$ as a Gaussian distribution with mean $(1-t)x_{0}+tx_{1}$ and a small constant standard deviation $\sigma$. This path interpolates between a noise sample $x_{0}\sim\mathcal{N}(0,I)$ and a data sample $x_{1}$. The corresponding target vector field is their difference, $u_{t}=x_{1}-x_{0}$. This formulation enables direct, simulation-free training of the model $v_{\theta}$ by regressing it against this target field.
@@ 3.2.1 Frame-wise Autoencoder
To learn a compact latent space, we train a VAE on individual radar frames. The architecture, inspired by esser2021taming, uses a hierarchical encoder $\mathcal{E}$ and decoder $\mathcal{D}$ with residual and self-attention blocks for high-fidelity reconstructions. The VAE is trained with a combination of a L1 reconstruction loss, a KL-divergence regularizer, and a PatchGAN adversarial loss (isola2017image) to enhance perceptual quality. After training, the VAE’s weights are frozen and it is used to encode inputs and decode latent predictions.
@@ 3.2.2 FlowCast Architecture
We propose FlowCast, which consists of the adaptation of Earthformer-UNet (gao2024prediff) for the CFM objective. FlowCast employs a U-Net-like encoder-decoder structure where the core building blocks are Cuboid Attention layers from Earthformer (gao2022earthformer). This mechanism efficiently processes spatiotemporal data by applying self-attention locally within 3D ”cuboids” of the data, capturing local dynamics, while global information is shared across the hierarchical U-Net structure. The model is conditioned on the flow time $t$, which is converted into an embedding and injected at each level of the network, enabling the model to accurately approximate the time-dependent vector field $v_{\theta}$. The architecture is illustrated in Figure 1.
@@
To generate an ensemble of forecasts, we solve the learned ODE starting from noise, using the Euler method (Hairer1993-za). This process is described in Algorithm 2.
@@ 4.1.1 Datasets
We evaluate on two 5-minute, 1 km-resolution radar datasets: SEVIR, a US benchmark, and ARSO, a Slovenian composite for a local deployment setting. For both, we predict 12 frames (1 hour) from 13 past frames (65 minutes), per the SEVIR Nowcasting Challenge protocol (veillette2020sevir).
@@
SEVIR (veillette2020sevir) provides over 10,000 weather events in a $\times384$ km US domain, each spanning 4 hours at 5-minute resolution. We use the 1-km Vertically Integrated Liquid (VIL) field. Following the standard chronological split, we extract 25-frame sequences (13 context, 12 target) with a stride of 12, yielding 36,351 training, 9,450 validation, and 12,420 test samples.
@@ 4.1.2 Evaluation
Following prior work (veillette2020sevir; gao2024prediff; gong2024cascast), we evaluate forecasts by converting radar fields to binary masks at given thresholds and computing the False Alarm Ratio (FAR), Critical Success Index (CSI), and Heidke Skill Score (HSS). For spatial validation, we compute the max-pooled CSI and Fractions Skill Score (FSS) over $16\times 16$ km neighborhoods (CSI-M-P16 and FSS-M-P16). We report the mean of these scores across all thresholds (”-M”) to evaluate general performance. Furthermore, to rigorously assess the detection of extreme weather events, we separately report the categorical metrics specifically at the highest intensity thresholds for each dataset.
@@ 4.1.3 Training Details
We train a separate Variational Autoencoder (VAE) for each dataset to create a specialized latent space. We follow the architecture and training procedure from rombach2022high, with a Kullback-Leibler divergence loss weight of 1e-4, the AdamW optimizer with a learning rate of 1e-4, and a batch size of 12. The compressed latent space dimensions are shown in Table 2.
@@
We train our CFM model for 200 epochs using the AdamW optimizer with a learning rate of 5e-4 and a cosine scheduler. We set the standard deviation of the I-CFM probability path to a small constant $\sigma=0.01$ (tong2024improving). We observed that the training process was notably stable; unlike diffusion objectives which often require complex loss weighting schedules, the I-CFM objective utilizes a simple regression loss that converged robustly without extensive hyperparameter tuning. Model checkpoints are maintained using an exponential moving average of weights (ho2020denoising), with a decay factor of 0.999, and we keep the model checkpoint with the highest CSI-M evaluated on a subset of the validation set. The model is trained with 4 NVIDIA H100 for 7 days, with a global batch size of 12. Further implementation details are provided in Appendix A.1.
@@ 4.1.4 Inference Details
Generating a forecast with FlowCast involves solving the learned ODE to transform a noise-initialized latent sequence into a prediction, conditioned on encoded past observations. Following the procedure outlined in 3.2.2, we use the Euler method (Hairer1993-za) with 10 steps as the ODE solver. To generate a probabilistic ensemble forecast, this process is repeated eight times with different initial noise samples $Z(0)\sim\mathcal{N}(0,\mathbf{I})$.
@@ 4.2 Comparison to the State of the Art
We evaluate FlowCast against four deterministic baselines: U-Net (veillette2020sevir), Earthformer (gao2022earthformer), Earthfarseer (wu2024earthfarsser), and SimVPv2 (tan2025simvpv2), as well as two probabilistic baselines: PreDiff (gao2024prediff) and CasCast (gong2024cascast). All models are trained following their publicly released code, with the training budget fixed at 200 epochs. For probabilistic models, we adopt our evaluation protocol by selecting the checkpoint with the highest CSI-M on a validation subset, using exponential moving average weights.
@@ 4.3.1 CFM objective against Diffusion
To isolate the benefits of the CFM objective, we compare FlowCast against a strong baseline using the same backbone architecture but trained with a diffusion objective. We trained a DDPM (ho2020denoising) for 1000 timesteps. For efficient inference, we employed a DDIM sampler (song2021denoising) with a varying number of steps. This provides a strong and practical baseline to evaluate FlowCast against a highly optimized diffusion process on the same powerful architecture.

### v3

@@ Abstract
Radar-based precipitation nowcasting, the task of forecasting short-term precipitation fields from previous radar images, is a critical problem for flood risk management and decision-making. While deep learning has substantially advanced this field, two challenges remain fundamental: the uncertainty of atmospheric dynamics and the efficient modeling of high-dimensional data. Diffusion models have shown strong promise by producing sharp, reliable forecasts, but their iterative sampling process is computationally prohibitive for time-critical applications. We introduce FlowCast, the first end-to-end probabilistic model leveraging Conditional Flow Matching (CFM) as a direct noise-to-data generative framework for precipitation nowcasting. Unlike hybrid approaches, FlowCast learns a direct noise-to-data mapping in a compressed latent space, enabling rapid, high-fidelity sample generation. Our experiments demonstrate that FlowCast establishes a new state-of-the-art in probabilistic performance while also exceeding deterministic baselines in predictive accuracy. A direct comparison further reveals the CFM objective is both more accurate and significantly more efficient than a diffusion objective on the same architecture, maintaining high performance with significantly fewer sampling steps. This work positions CFM as a powerful and practical alternative for high-dimensional spatiotemporal forecasting.  [^1] Our implementation is available at https://github.com/b-rbmp/FlowCast.
@@ 1 Introduction
Accurate and timely short-term precipitation forecasts, or nowcasting, are of paramount importance due to their significant socio-economic impacts, such as issuing flood warnings and managing water resources. Precipitation nowcasting, as defined in this work, involves predicting a sequence of future radar images from historical observations for the immediate future up to a few hours (An et al., 2025). Traditional methods, like Eulerian and Lagrangian persistence (Germann and Zawadzki, 2002), rely on advecting the current precipitation field. However, their simplified physical assumptions limit their ability to capture the complex, non-linear dynamics of atmospheric processes, especially for rapidly evolving weather systems (Prudden et al., 2020).
Deep learning has introduced a paradigm shift in precipitation nowcasting. Deterministic models based on recurrent and transformer architectures learn complex spatiotemporal patterns directly from large volumes of radar data (Prudden et al., 2020; An et al., 2025). While these models outperform traditional methods, optimizing for metrics like Mean Squared Error (MSE) compels them to produce a single, best-guess forecast. This often results in overly smooth predictions at longer lead times, failing to capture the inherent uncertainty in precipitation evolution and underrepresenting high-impact weather events.
To address this, probabilistic generative models have become central to modern nowcasting, aiming to predict a distribution over many plausible futures. Diffusion models (Ho et al., 2020), in particular, have emerged as the state-of-the-art, producing sharp and reliable ensemble forecasts (Gao et al., 2023; Leinonen et al., 2023; Gong et al., 2024). However, their reliance on an iterative denoising process, often requiring hundreds of function evaluations for a single forecast, makes them computationally expensive. This high Number of Function Evaluations (NFE) poses a significant barrier in time-critical scenarios where rapid ensemble generation is crucial.
This work introduces FlowCast, a novel probabilistic nowcasting model built on Conditional Flow Matching (CFM) (Lipman et al., 2023; Tong et al., 2024), a powerful and efficient alternative designed for rapid sampling. While recent work has applied rectified flows for the deterministic refinement of blurry forecasts (Feng et al., 2025), FlowCast is, to our knowledge, the first to successfully apply CFM as a full, noise-to-data generative model for this task. We demonstrate that FlowCast alleviates the tension between accuracy and efficiency, establishing a new state-of-the-art by exceeding the performance of leading diffusion models while offering a superior performance-cost trade-off.
@@
  • We establish a new state-of-the-art in both probabilistic performance and predictive accuracy on two diverse radar datasets, the benchmark SEVIR dataset (Veillette et al., 2020) and the local ARSO dataset.
@@ 2.1 Deterministic Nowcasting
Deep learning for precipitation nowcasting has evolved from RNN-based architectures to Transformer-based models. Early work includes ConvLSTM (Shi et al., 2015), extending LSTMs with convolutions for spatiotemporal data, and the PredRNN family (Wang et al., 2017; 2023), which introduced a spatiotemporal memory flow for improved long-range dependency modeling. More recently, Transformer architectures like Earthformer (Gao et al., 2022) and Earthfarseer (Wu et al., 2024) have set new benchmarks by using attention to model complex global dynamics. A common limitation of deterministic models is that they produce overly smooth forecasts when trained with pixel-wise losses (e.g., MSE), as they average over possible futures.
@@ 2.2 Probabilistic Nowcasting
GANs (Ravuri et al., 2021) were an early approach for producing sharp forecasts but suffer from training instability. Diffusion models (Ho et al., 2020) have recently emerged as the state-of-the-art, offering stable training and high-quality samples. PreDiff (Gao et al., 2023) and LDCast (Leinonen et al., 2023) are prominent latent diffusion models for ensemble forecasting. A notable hybrid is CasCast (Gong et al., 2024), which uses a deterministic model for large-scale patterns and a conditional diffusion model to refine stochastic details.
@@
Generative modeling with flows offers an attractive alternative to diffusion. Traditional Continuous Normalizing Flows (CNFs) (Chen et al., 2018) model data via ODEs but require expensive numerical integration during training to compute likelihoods, making them computationally prohibitive for high-dimensional spatiotemporal data. Conditional Flow Matching (CFM) (Lipman et al., 2023) and Rectified Flows (Liu et al., 2023) overcome this by regressing a vector field against a conditional probability path, enabling simulation-free training. However, while standard Rectified Flows typically utilize a singular conditional path (effectively $\sigma\to 0$), our application of Independent CFM (I-CFM) incorporates a Gaussian probability path with $\sigma>0$. This ”thickens” the training trajectory, providing crucial regularization that stabilizes the learning of the vector field for high-dimensional data compared to the singular paths of rectified flows.
Crucially, while diffusion models rely on stochastic denoising paths that are often curved and require many sampling steps, CFM allows for learning straight-line ODE trajectories between noise and data (Tong et al., 2024). This geometric property enforces a direct mapping that preserves temporal coherence and allows for rapid sampling. While Feng et al. (2025) recently used a rectified flow module to strictly refine deterministic forecasts, FlowCast applies CFM as a standalone probabilistic generative model. This allows it to learn the full noise-to-data distribution and capture multimodal uncertainty without relying on a deterministic base forecast.
@@ 3.2 Latent Conditional Flow Matching
To reduce the high computational cost of generative modeling, we adopt a two-stage approach inspired by latent diffusion models (Rombach et al., 2022). A Variational Autoencoder (VAE) (Kingma and Welling, 2014) compresses high-dimensional radar frames into low-dimensional latents, which are used to train a generative model in the latent space.
Our generative model is built on the Conditional Flow Matching (CFM) framework (Lipman et al., 2023), which trains a continuous normalizing flow by learning a vector field $v_{\theta}$ that maps samples from a prior distribution (e.g., Gaussian) to the target data distribution. We use Independent CFM (I-CFM) (Tong et al., 2024), which defines a probability path $p_{t}(x_{t}|x_{0},x_{1})$ as a Gaussian distribution with mean $(1-t)x_{0}+tx_{1}$ and a small constant standard deviation $\sigma$. This path interpolates between a noise sample $x_{0}\sim\mathcal{N}(0,I)$ and a data sample $x_{1}$. The corresponding target vector field is their difference, $u_{t}=x_{1}-x_{0}$. This formulation enables direct, simulation-free training of the model $v_{\theta}$ by regressing it against this target field.
@@ 3.2.1 Frame-wise Autoencoder
To learn a compact latent space, we train a VAE on individual radar frames. The architecture, inspired by Esser et al. (2021), uses a hierarchical encoder $\mathcal{E}$ and decoder $\mathcal{D}$ with residual and self-attention blocks for high-fidelity reconstructions. The VAE is trained with a combination of a L1 reconstruction loss, a KL-divergence regularizer, and a PatchGAN adversarial loss (Isola et al., 2017) to enhance perceptual quality. After training, the VAE’s weights are frozen and it is used to encode inputs and decode latent predictions.
@@ 3.2.2 FlowCast Architecture
We propose FlowCast, which consists of the adaptation of Earthformer-UNet (Gao et al., 2023) for the CFM objective. FlowCast employs a U-Net-like encoder-decoder structure where the core building blocks are Cuboid Attention layers from Earthformer (Gao et al., 2022). This mechanism efficiently processes spatiotemporal data by applying self-attention locally within 3D ”cuboids” of the data, capturing local dynamics, while global information is shared across the hierarchical U-Net structure. The model is conditioned on the flow time $t$, which is converted into an embedding and injected at each level of the network, enabling the model to accurately approximate the time-dependent vector field $v_{\theta}$. The architecture is illustrated in Figure 1.
@@
To generate an ensemble of forecasts, we solve the learned ODE starting from noise, using the Euler method (Hairer et al., 1993). This process is described in Algorithm 2.
@@ 4.1.1 Datasets
We evaluate on two 5-minute, 1 km-resolution radar datasets: SEVIR, a US benchmark, and ARSO, a Slovenian composite for a local deployment setting. For both, we predict 12 frames (1 hour) from 13 past frames (65 minutes), per the SEVIR Nowcasting Challenge protocol (Veillette et al., 2020).
@@
SEVIR (Veillette et al., 2020) provides over 10,000 weather events in a $\times384$ km US domain, each spanning 4 hours at 5-minute resolution. We use the 1-km Vertically Integrated Liquid (VIL) field. Following the standard chronological split, we extract 25-frame sequences (13 context, 12 target) with a stride of 12, yielding 36,351 training, 9,450 validation, and 12,420 test samples.
@@ 4.1.2 Evaluation
Following prior work (Veillette et al., 2020; Gao et al., 2023; Gong et al., 2024), we evaluate forecasts by converting radar fields to binary masks at given thresholds and computing the False Alarm Ratio (FAR), Critical Success Index (CSI), and Heidke Skill Score (HSS). For spatial validation, we compute the max-pooled CSI and Fractions Skill Score (FSS) over $16\times 16$ km neighborhoods (CSI-M-P16 and FSS-M-P16). We report the mean of these scores across all thresholds (”-M”) to evaluate general performance. Furthermore, to rigorously assess the detection of extreme weather events, we separately report the categorical metrics specifically at the highest intensity thresholds for each dataset.
@@ 4.1.3 Training Details
We train a separate Variational Autoencoder (VAE) for each dataset to create a specialized latent space. We follow the architecture and training procedure from Rombach et al. (2022), with a Kullback-Leibler divergence loss weight of 1e-4, the AdamW optimizer with a learning rate of 1e-4, and a batch size of 12. The compressed latent space dimensions are shown in Table 2.
@@
We train our CFM model for 200 epochs using the AdamW optimizer with a learning rate of 5e-4 and a cosine scheduler. We set the standard deviation of the I-CFM probability path to a small constant $\sigma=0.01$ (Tong et al., 2024). We observed that the training process was notably stable; unlike diffusion objectives which often require complex loss weighting schedules, the I-CFM objective utilizes a simple regression loss that converged robustly without extensive hyperparameter tuning. Model checkpoints are maintained using an exponential moving average of weights (Ho et al., 2020), with a decay factor of 0.999, and we keep the model checkpoint with the highest CSI-M evaluated on a subset of the validation set. The model is trained with 4 NVIDIA H100 for 7 days, with a global batch size of 12. Further implementation details are provided in Appendix A.1.
@@ 4.1.4 Inference Details
Generating a forecast with FlowCast involves solving the learned ODE to transform a noise-initialized latent sequence into a prediction, conditioned on encoded past observations. Following the procedure outlined in 3.2.2, we use the Euler method (Hairer et al., 1993) with 10 steps as the ODE solver. To generate a probabilistic ensemble forecast, this process is repeated eight times with different initial noise samples $Z(0)\sim\mathcal{N}(0,\mathbf{I})$.
@@ 4.2 Comparison to the State of the Art
We evaluate FlowCast against four deterministic baselines: U-Net (Veillette et al., 2020), Earthformer (Gao et al., 2022), Earthfarseer (Wu et al., 2024), and SimVPv2 (Tan et al., 2025), as well as two probabilistic baselines: PreDiff (Gao et al., 2023) and CasCast (Gong et al., 2024). All models are trained following their publicly released code, with the training budget fixed at 200 epochs. For probabilistic models, we adopt our evaluation protocol by selecting the checkpoint with the highest CSI-M on a validation subset, using exponential moving average weights.
@@ 4.3.1 CFM objective against Diffusion
To isolate the benefits of the CFM objective, we compare FlowCast against a strong baseline using the same backbone architecture but trained with a diffusion objective. We trained a DDPM (Ho et al., 2020) for 1000 timesteps. For efficient inference, we employed a DDIM sampler (Song et al., 2021) with a varying number of steps. This provides a strong and practical baseline to evaluate FlowCast against a highly optimized diffusion process on the same powerful architecture.
@@ 5 Conclusion
Acknowledgments
This work was supported by the Slovenian Research and Innovation Agency (ARIS) research core funding P2-0209 (Jana Faganeli Pucer). Special thanks go to Janko Merše and Dr. Matic Šavli from the Slovenian Environment Agency (ARSO) for their support and for providing the ARSO dataset.